### v2

@@ 3.1. Task Formulation
Formally, given a financial document $D$ consisting of multiple textual contents and structured tables, a set of predefined value data types $L$ specified by the XBRL specification, and a taxonomy database $\mathcal{T}$ containing a set of financial semantic concepts, the XBRL tagging task is to identify all relevant numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. Here, Tag denotes the linked concept in $\mathcal{T}$ (e.g. “us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItem, and integerItemType). Each numerical fact is assigned a single, most specific tag from the taxonomy, without overlapping. Specifically, monetaryItemType refers to a financial amount, such as revenue or expenses, typically reported in currency units (e.g., USD). percentItemType represents a rate or ratio, expressed as a decimal (e.g., 25%). sharesItemType indicates the number of shares, such as stocks held or issued. perShareItemType captures values reported on a per-share basis, like earnings per share. Lastly, integerItemType is used for whole number counts, such as the number of employees or transactions.
@@
(1) $f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$
@@
Inspired by information extraction and alignment works (Wu et al., 2019; Wadden et al., 2019; Mukherjee et al., 2022; Wang et al., 2022) and after discussed with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type. Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
@@
(2) $f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$
@@
(3) $f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$
where $e$ is a numerical entity identified in the document $D=(S,T)$, $l\in L$ is its predicted data type from FinNI, $C_{e}$ denotes the contextual information surrounding $e$ in $D$, and $\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$ is a financial taxonomy containing $n$ uniquely defined and semantically grounded concepts. The model is required to assign a concept $\hat{c}\in\mathcal{T}$ that best reflects the meaning of $e$ in its structural context.
3.2. Raw Data Collection
We compiled 142 annual 10-K reports filed in 2023 and 2024 by publicly listed companies from the SEC [^6] https://www.sec.gov/, as shown in Table 2. To ensure diversity, the collection covers all 11 major industry sectors, with about 95% of companies distributed across 31 U.S. states and the remaining 5% based outside the United States. Using BeautifulSoup, we parsed these reports and extracted 319,893 narrative sentences (roughly 76 million characters) along with 21,576 financial tables. Company-level details are provided in Appendix D.2. Each report follows the SEC’s XBRL filing standard, meaning that the underlying instance, schema, and linkbase documents already contain machine-readable tags (ground-truth tags) that link textual and tabular values to US-GAAP taxonomy concepts. We included all sections that may reference or embed XBRL-tagged content to ensure comprehensive coverage. For the justification of our benchmark granularity, please see the Appendix E.
Table 2. Statistics of collected financial reports.
@@ 3.2. Raw Data Collection
3.3. Data Curation
To construct the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark and ensure both efficiency and fairness in evaluation, we designed a two-stage filtering and annotation pipeline. This pipeline systematically processes the raw 10-K filings to retain representative, high-quality instances that collectively cover all 142 companies, five valid entity types, and every numerical US-GAAP concept observed in the corpus. The overall workflow is illustrated in Algorithms 1 and 2.
Stage 1: Filtering Raw Instances (Algorithm 1)
The first stage aims to remove noisy or redundant samples while ensuring comprehensive coverage of both entity types and US-GAAP concepts. Each narrative sentence or table instance is automatically examined by a rule-based parser that checks if it contains numeric entities tagged with valid US-GAAP concepts. The parser performs schema-level consistency checks based on the taxonomy definitions to ensure tagging correctness. Instances without valid entities are classified as NEG (negative). Those containing valid entities are treated as positive and further divided into two categories: (1) POS_SELECT, which introduces at least one new US-GAAP concept not yet covered; and (2) POS_CANDIDATE, which contains only previously covered concepts. This design ensures that each concept appears at least once in the dataset while minimizing redundancy across companies and reports.
Figure 2. Statistics of numerical entity types.
To determine which entity types should be considered valid, we first profiled all XBRL numeric item types that appeared in the raw corpus and summarized the results in Figure 2. For each type, we calculated both its cross-company coverage (the number of companies in which it occurs) and its overall frequency (the proportion of all annotated numeric entities belonging to that type). The figure shows a clear long-tailed distribution: a total of eleven numeric item types were identified, but only a few dominate in both coverage and frequency. The five types that consistently appear across nearly all companies and collectively account for the majority of entities are monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType. These five types form the label space for all subsequent stages. Under this definition, Algorithm 1 treats an instance as a potential positive only if it contains at least one US-GAAP–tagged entity belonging to this label space; otherwise it is assigned to NEG.
To maintain structural fidelity, additional validity rules are applied. For narrative texts, we retain only segments longer than 20 characters that contain numerical tokens. For tables, we preserve the <table>, <tr>, <td>, and <th> tags to ensure layout integrity, while tables with headers only (i.e., without numeric values) are excluded.
Input: 10-K reports; US-GAAP taxonomy; 5 valid entity types
Output: POS_SELECT, POS_CANDIDATE, NEG
Initialize: an empty set of covered concepts; empty lists for POS_SELECT, POS_CANDIDATE, and NEG;
foreach instance (texts or tables) in 10-K reports do
  Check if the instance contains numeric entities that are both: (i) tagged with a US-GAAP concept, and (ii) of a valid entity type;
 if no such entities are found then
     Add the instance to NEG;
 else
     Identify all US-GAAP concepts in this instance;
     if the instance introduces at least one new concept not yet covered then
        Add the instance to POS_SELECT and update the covered concepts set;
    else
        Add the instance to POS_CANDIDATE;
return {POS_SELECT, POS_CANDIDATE, NEG};
Algorithm 1 Screening of Text and Table Instances
Stage 2: Company-Capped Sampling and Minimal Cover (Algorithm 2)
The second stage refines the positive pool by enforcing both per-company balance and concept coverage completeness. For each modality (text or table), we first perform company-capped seeding: from each company’s POS_SELECT, up to $K=10$ positive instances are retained as the seed set, while the remaining positives are moved into a candidate pool. We then mark all US-GAAP concepts covered by the seed and identify those still uncovered. Finally, a greedy minimal-cover procedure is executed: candidates are iteratively selected according to the number of uncovered concepts they can cover, with ties broken by shorter text length, fewer entities, or earlier order. Each selected candidate is added to the final set $P^{*}$, and the process continues until all concepts are covered or the candidate pool is exhausted. The resulting positive set $P^{*}$ thus maintains full US-GAAP concept coverage, balanced company representation, and reduced redundancy, forming a compact yet coverage-complete dataset.
Input: POS_SELECT grouped by company (run separately for text and table) from algorithm 1;
   All concepts observed in the POS_SELECT;
   Cap $K=10$ (maximum positives per company for seeding)
Output: Final positive instances $P^{*}$ and leftover instances
Seed selection. For each company, keep up to $K$ positives as the seed set. Put the rest into a new candidate pool.
Check coverage. Mark all concepts already covered by the seed. Identify which concepts are still uncovered.
Greedy covering. While uncovered concepts remain:
  1. (1)
    For each candidate, count how many uncovered concepts it can cover.
  2. (2)
    Pick the candidate with the highest count (ties broken by shorter text, fewer entities, or earlier order).
  3. (3)
    Add the candidate to the final positives, update the covered concepts, and remove it from the pool.
Finalize. The final positives $P^{*}$ are the union of the seed and the greedy additions. The leftover candidates form the candidate set.
Algorithm 2 Company-Capped Seeding + Minimal Cover (for Text and Table)
Final Dataset and Validation
After two-stage filtering, we obtained 3,084 textual sentences and 3,067 table sequences from financial filings (Table 3). Among them, 2,054 text and 2,045 table instances are positives containing valid XBRL tagging information, while 1,030 text and 1,022 table instances are negatives. Textual inputs contain on average $96.25$ tokens, whereas table sequences are much longer ($1,204.72$ tokens on average). Overall, the dataset includes 62,668 annotated entities linked to 3,953 unique US-GAAP concepts across the five valid entity types.
Based on this foundation, we further derive two subtask-specific datasets: FinNI-eval, targeting numerical entity identification, and FinCL-eval, targeting concept linking. Both are the first datasets in the financial domain to emphasize structured information extraction and concept-level alignment. Their detailed statistics are reported in Table 4.
To verify annotation reliability, two domain experts independently reviewed 100 randomly sampled entries. The results indicate a 96% raw agreement and a Cohen’s Kappa of 81.13, confirming the robustness and reproducibility of our semi-automated annotation pipeline. The validation guideline is provided in Appendix F, and a complete unfiltered version of the benchmark is included in Appendix D.1.
@@ 3.3. Data Curation
Table 3. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
@@
Positive instances 2,054 2,045
Negative instances 1,030 1,022
Avg. Tokens/S $96.25\pm 72.75$ $1,204.72\pm 1,162.41$
Avg. Entities/S $1.88\pm 2.11$ $18.54\pm 29.14$
Avg. Concepts/S $1.88\pm 2.11$ $18.54\pm 29.14$
Total Entities 62,668
@@
Table 4. Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
@@
3.4. FinNI-eval Dataset Construction
Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 6,151 instances derived from the annotated sentences and tables. Specifically, the dataset comprises two components: the input block and the answer block. The input block includes a task instruction and the input content. The instruction defines the FinNI task by specifying the responsibilities of the LLMs, the definitions of entity types, and the rules to be followed. The answer block is formatted as a list aligned with the input block. The answer is a JSON list of all identified entity values and types or an empty JSON list. The specific prompt template is shown in Figure 9 in the Appendix.
@@ 3.4. FinNI-eval Dataset Construction
Answer: {"results": [{"value": <numeric entity>, "type": <entity type>}]}
        or {"results": []}
3.5. FinCL-eval Dataset Construction
Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 62,668 query-answer pairs for numerical entity normalization. In addition, we built a US-GAAP taxonomy database containing 17,688 unique financial concepts. To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3. Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
@@ 3.5. FinCL-eval Dataset Construction
3.6. Evaluation
3.6.1. Evaluation Framework
@@
We propose a unified evaluation framework for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark to assess LLM performance. As shown in Figure 3, given a financial report $D$, the framework reformulates tagging into two subtasks, FinNI and FinCL, to generate structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. This design jointly evaluates an LLM’s zero-shot ability in fact extraction and concept alignment. For fair comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 (Sokolova and Lapalme, 2009). FinNI is evaluated using pair-level metrics, while FinCL uses accuracy. Full metric definitions are provided in Appendix G.
Figure 3. A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. Note: The Sentence is drawn from financial report $D$, and the Query provides an example that incorporates the entity 4.9. The black arrows show the overall framework flow, green arrows denote the FinNI process, red arrows indicate the FinCL process, red bidirectional arrows represent retrieval interactions with the taxonomy, and red dashed arrows mark the recomposed input between the query and retrieved candidates.
FinNI Evaluation: We use pair-level metrics to evaluate the extraction performance on FinNI. The objective is to evaluate the LLM’s ability to distinguish between structured and unstructured contexts and to extract financial values based on its semantic understanding and domain-specific knowledge.
@@
FinCL Evaluation: To assess whether LLMs can perform fine-grained XBRL entity normalization, we reformulate the task as a retrieval-reranking problem, where LLMs are used to disambiguate and select the most appropriate taxonomy concept (reranking stage) from a reduced candidate set (retrieval stage), avoiding the impracticality of direct multi-thousand-way classification (prompt template used for reranking in the FinCL subtask is shown in Figure 10 in Appendix).
@@
To this end, we first generate embeddings for each taxonomy concept using text-embedding-3-small [^7] https://platform.openai.com/docs/models/text-embedding-3-small, and retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ based on semantic similarity. Therefore, the LLMs are required to rerank the retrieved candidates by leveraging deeper contextual understanding to select the best-matched tag $\hat{c}\in\mathcal{T}_{h}$, formalized as:
(4) $f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$
3.6.2. Evaluation Models

### v3

@@ 3.1 Task Formulation
Formally, given a financial document $D$ consisting of multiple textual contents and structured tables, a set of predefined value data types $L$ specified by the XBRL specification, and a taxonomy database $\mathcal{T}$ containing a set of financial semantic concepts, the XBRL tagging task is to identify all relevant numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. Here, Tag denotes the linked concept in $\mathcal{T}$ (e.g. “us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType). Each numerical fact is assigned a single, most specific tag from the taxonomy, without overlapping. Specifically, monetaryItemType refers to a financial amount, such as revenue or expenses, typically reported in currency units (e.g., USD). percentItemType represents a rate or ratio, expressed as a decimal (e.g., 25%). sharesItemType indicates the number of shares, such as stocks held or issued. perShareItemType captures values reported on a per-share basis, like earnings per share. Lastly, integerItemType is used for whole number counts, such as the number of employees or transactions.
@@
$f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$ (1)
@@
Inspired by information extraction and alignment works wu2019scalable; wadden2019entity; mukherjee2022ectsum; wang2022conditional and after discussing with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type. Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
@@
$f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$ (2)
@@
$f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$ (3)
where $e$ is a numerical entity extracted from $D=(S,T)$, and $l\in L$ is its associated data type provided by the FinNI stage. $C_{e}$ represents the structural context of $e$, which encompasses linguistic dependencies for narrative mentions and hierarchical row-column coordinates for tabular mentions. $\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$ denotes the full-scope US-GAAP taxonomy containing $n$ uniquely defined and semantically grounded concepts. The objective of FinCL is to identify the optimal concept $\hat{c}\in\mathcal{T}$ that maximizes the semantic alignment between the fact $e$ and its specific context $C_{e}$ within the high-dimensional taxonomy space.
3.2 Raw Data Collection
We compiled 142 annual 10-K filings from publicly listed U.S. companies submitted in 2023–2024 (Table 2). To ensure coverage diversity, the collection spans all 11 major industry sectors, with about 95% of companies located across 31 U.S. states and the remainder based internationally. Using BeautifulSoup, we parsed the reports and extracted 319,893 narrative sentences and 21,576 financial tables. Company-level statistics are provided in Appendix C.1. All filings follow the SEC Inline XBRL (iXBRL) format, which embeds machine-readable tags inside HTML disclosures and links textual and tabular values to US-GAAP taxonomy concepts through schema and linkbase definitions. However, these regulatory tags are designed for compliance rather than model evaluation. They provide little coverage of negative cases and contain substantial structural noise. We therefore construct a high-fidelity evaluation suite that repurposes this raw regulatory metadata into a standardized benchmark for financial fact extraction and concept linking. Additional discussion of our benchmark granularity is provided in Appendix C.2.
Table 2: Statistics of collected financial reports.
@@ 3.2 Raw Data Collection
3.3 Large-Scale Automated Annotation
Our automated pipeline produces a large, high-fidelity corpus reflecting real financial disclosures. As summarized in Table 3, the dataset contains 15,986 narrative sentences and 12,801 tables, yielding 261,457 numerical entities linked to 3,953 unique US-GAAP concepts. Including both valid facts and naturally occurring noise, it supports a realistic evaluation of LLMs under regulatory-style conditions.
@@ 3.3 Large-Scale Automated Annotation
Table 3: Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
@@
Positive instances 7,768 8,709
Negative instances 8,218 4,092
Avg. Tokens/S $80.91\pm 63.62$ $1212.42\pm 1421.76$
Avg. Entities/S $1.24\pm 1.82$ $18.87\pm 37.16$
Avg. Concepts/S $1.24\pm 1.82$ $18.87\pm 37.16$
Total Entities 261,457
@@
Automated Annotation Engine. We build a rule-based parser (Algorithm 1, Appendix C.3) that reconstructs semantic links between human-readable disclosures and Inline XBRL (iXBRL) metadata. For each instance $i$, the parser locates ix:nonfraction elements and extracts the value $v$, concept $c$, and data type $\tau$, while restricting extraction to atomic leaf elements $\epsilon$ to avoid duplicates from nested tags.
Valid Entity Types. To define the label space, we profile all numeric XBRL item types. As shown in Figure 2, the distribution is long-tailed: although eleven types appear, five dominate. We therefore restrict $\mathcal{L}$ to monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType. An instance is labeled POSITIVE ($\mathcal{P}$) only if it contains at least one concept $c\in\mathcal{T}$ from $\mathcal{L}$; otherwise it is labeled NEGATIVE ($\mathcal{N}$).
Structural Refinement and Filtering. For reliable evaluation, we keep only text segments longer than 20 characters and deduplicate content globally using $\mathcal{S}_{text}$. For tables, we retain structural tags (<table>, <tr>, <td>, <th>) but discard tables without numerical facts. This filtering preserves semantic accuracy while maintaining realistic layout structure.
Figure 2: Statistics of numerical entity types.
3.4 Data Quality Control
Validating all 261,457 entities is infeasible. To obtain a rigorous quality upper bound, we adopt a strategic sampling scheme based on a greedy minimal-cover algorithm (Algorithm 2, Appendix D.1). Rather than random sampling, we construct a diversity-maximized subset $\mathcal{P}^{*}$ that satisfies two goals: (1) Company Balance, limiting bias from individual reporting styles, and (2) Taxonomy Completeness, ensuring that each US-GAAP concept observed in the corpus appears at least once. We seed $\mathcal{P}^{*}$ with capped samples per company, then iteratively add instances that cover the largest number of uncovered concepts, producing a compact but representative audit set. The final subset contains 928 positive sentence instances and 1,021 positive table instances, serving as a taxonomy-wide stress test reviewed by two junior analysts and one senior adjudicator (the demography of these annotators is shown in Appendix D.2).
Review Protocol. The audit followed two stages. A pilot phase aligned interpretations of the Verification Guideline (Appendix D.3), after which analysts independently verified whether extracted triplets $(e,\tau,c)$ were fully correct. The senior expert adjudicated disagreements to create the final gold set.
Agreement Results. On the independent round, annotators achieved 96% raw agreement and Cohen’s $\kappa=0.81$, indicating substantial agreement landis1977measurement, the equation is shown in Appendix D.4. This high-verified accuracy on the most semantically diverse subset provides strong evidence for the reliability of our automated annotation pipeline.
3.5 Task-Specific Dataset Construction
Following validation, we integrated the corrected $\mathcal{P}^{*}$ subset back into the full corpus to form a high-fidelity hybrid dataset. This refined pool ensures that every US-GAAP concept in the taxonomy has at least one expert-verified representation. From this foundation, we derive two subtask-specific datasets: FinNI-eval, targeting numerical entity identification, and FinCL-eval, targeting concept linking. Both are the first datasets in the financial domain to emphasize structured information extraction and concept-level alignment. Their detailed statistics are reported in Table 4.
Table 4: Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
@@ 3.5 Task-Specific Dataset Construction
3.5.1 FinNI-eval Dataset Construction
Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 28,787 instances derived from the annotated sentences and tables. Specifically, the dataset comprises two components: the input block and the answer block. The input block includes a task instruction and the input content. The instruction defines the FinNI task by specifying the responsibilities of the LLMs, the definitions of entity types, and the rules to be followed. The answer block is formatted as a list aligned with the input block. The answer is a JSON list of all identified entity values and types or an empty JSON list. The specific prompt template is shown in Figure 9.
@@ 3.5.1 FinNI-eval Dataset Construction
Answer: {"results": [{"Fact": <numeric entity>,
        "Type": <entity type>}]} or {"results": []}
3.5.2 FinCL-eval Dataset Construction
Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 261,457 query-answer pairs for numerical entity normalization. In addition, we built a US-GAAP taxonomy database containing 17,688 unique financial concepts. To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3. Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
@@ 3.5.2 FinCL-eval Dataset Construction
3.6 Evaluation
3.6.1 Evaluation Framework
@@
We propose a unified evaluation framework for 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. As illustrated in Figure 3, each input is decomposed into two subtasks, FinNI and FinCL, which together produce structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. This design jointly evaluates zero-shot fact extraction and concept alignment. For comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 sokolova2009systematic. FinNI is evaluated using pair-level metrics, and FinCL with accuracy. Full metric definitions are in Appendix E.
Figure 3: A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. Note: The Input content is drawn from the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset, and the “Shared LLMs” are used for identification in the FinNI task and reranking in the FinCL task.
FinNI Evaluation. We adopt pair-level metrics to measure extraction quality, focusing on whether models can locate financial values from mixed text–table contexts based on semantic cues.
@@
FinCL Evaluation. To test fine-grained numeric entity normalization, we formulate FinCL as a retrieval–reranking task: a retrieval stage produces a candidate set of taxonomy concepts, and the LLM then reranks candidates to select the most appropriate tag, avoiding impractical extreme classification (prompt template in Figure 10, Appendix).
@@
To this end, we embed each taxonomy concept using text-embedding-3-small [^5] https://platform.openai.com/docs/models/text-embedding-3-small , then retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ by semantic similarity. The LLM subsequently reranks these candidates using contextual reasoning to select the best tag $\hat{c}\in\mathcal{T}_{h}$:
$f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$ (4)
3.6.2 Evaluation Models